In the summer a large pool evaporates water at 15% per day. Given 25,700 gallons of water, which function models the pool’s loss?
f(x) = 25700(0.15)^x
f(x) = 25700(0.85)^x
f(x) =25700(1.85)^x
f(x) =25700(1.15)^x






Answer by Alan3354(69443)

In the summer a large pool evaporates water at 15% per day. Given 25,700 gallons of water, which function models the pool’s loss?
f(x) = 25700(0.15)^x
f(x) = 25700(0.85)^x
f(x) =25700(1.85)^x
f(x) =25700(1.15)^x
----------
The 1st one, f(x) = 25700(0.15)^x models the pool's loss for the 1st day only.
None of the others model its loss.
-----
The 2nd one, f(x) = 25700(0.85)^x models the amount remaining in the pool - not the loss.